# Matched Filtering-Based Channel Estimation for AFDM Systems in Doubly Selective Channels

### v1

@@ IV-C MF CE for FDFD channels
2:  According to (48), $\hat{\varUpsilon}_{i}$ and $\hat{\varUpsilon}_{i}^{\prime}$ are obtained.
3:  According to (50), $\hat{\gamma}_{i}$ and $\hat{\gamma}_{i}^{\prime}$ are estimated.
4:  According to (51) and (52), $\hat{l}_{\tau_{i}}^{\left(t\right)}$ and $\hat{k}_{\nu_{i}}^{\left(t\right)}$ are obtained.
5:  $\hat{h}^{\left(t\right)}$ can be obtained by (46).
6:  $\hat{\boldsymbol{l}}_{\tau}=\left[\hat{\boldsymbol{l}}_{\tau},\hat{l}_{\tau_{i}}^{\left(t\right)}\right]$, $\hat{\boldsymbol{k}}_{\nu}=\left[\hat{\boldsymbol{k}}_{\nu},\hat{k}_{\nu_{i}}^{\left(t\right)}\right]$, $\hat{\boldsymbol{h}}=\left[\hat{\boldsymbol{h}},\hat{h}^{\left(t\right)}\right]$.
7:  The $\boldsymbol{y}_{\mathrm{T}_{1}}^{\left(t+1\right)}$ and $\boldsymbol{y}_{\mathrm{T}_{2}}^{\left(t+1\right)}$ are obtained by (54).
8:  $t=t+1$.
9:  until $t=T_{\mathrm{iter}}$ or $\left|\small{\frac{\left\|\boldsymbol{y}_{\mathrm{T}}^{\left(t+1\right)}\right\|-\left\|\boldsymbol{y}_{\mathrm{T}}^{\left(t\right)}\right\|}{\left\|\boldsymbol{y}_{\mathrm{T}}^{\left(t\right)}\right\|}}\right|\leqslant\sigma$.
9:  $\left(\hat{\boldsymbol{l}}_{\tau},\hat{\boldsymbol{k}}_{\nu},\hat{\boldsymbol{h}}\right)$.
IV-D MF-GFS CE for FDFD channels
In Subsection IV-C, an MF-based CE scheme is proposed. This scheme eliminates the matrix inversion operation, decouples joint estimation, and narrows the search range to significantly reduce the computational complexity. However, the trade-off between the complexity and estimation performance of the MF CE scheme yet depends on the $\rho$ in (49). The larger the $\rho$, the finer the grid $\boldsymbol{\varGamma}$, enabling more accurate estimation of fractional part. In this case, $\left(\hat{\boldsymbol{l}}_{\tau},\hat{\boldsymbol{k}}_{\nu},\hat{\boldsymbol{h}}\right)$ will become more accurate. However, the cost of increasing $\rho$ is that the complexity will inevitably increase. To address this issue, we introduce a MF-GFS CE scheme based on the GFS algorithm in this subsection, by significantly reducing the search range of fractional part to reduce complexity and estimating $\left(\hat{\boldsymbol{l}}_{\tau},\hat{\boldsymbol{k}}_{\nu},\hat{\boldsymbol{h}}\right)$ more accurately.
GFS is an unconstrained nonlinear optimization method for unimodal functions [35, 36, 37]. The GFS algorithm unevenly divides the search interval into two parts through the ratio of two consecutive generalized Fibonacci numbers (GFNs) to achieve an efficient search. The $i$-th GFN is denoted by $S_{i}\left(a,b,p,q\right)$, where $i$ is a non-negative integer. Note that the conditions that must be satisfied to estimate the optimal value of a function using the GFS algorithm are:
  • $a+b>0$, and $p$ as well as $q$ are two positive integers.
  • The function to be optimized needs to be a bounded unimodal function.
@@ IV-D MF-GFS CE for FDFD channels
In (50), the objective function $\left|\boldsymbol{\mathcal{T}}^{\mathrm{H}}\left(\hat{\varUpsilon}_{i}+\gamma\right)\boldsymbol{y}_{\mathrm{T}}\right|^{2}$ is a bounded unimodal function.
@@
After dividing the search interval by two consecutive GFNs, the interval of uncertainty (IU) containing the optimal value ${\gamma}_{i}$ needs to be determined. For convenience, assume that the search interval is $\left[x_{s},x_{f}\right]$, and $x_{s}<x_{1}<x_{2}<x_{f}$. The new IU $\left[x_{s}^{\prime},x_{f}^{\prime}\right]$ can be expressed as
@@
Note that $\left|x_{f}^{\prime}-x_{s}^{\prime}\right|<\left|x_{f}-x_{s}\right|$.
@@
The process of the GFS algorithm for estimating $\hat{\gamma}_{i}$ is shown in Fig. 7. Note that it is necessary to set a maximum GFN $S_{N_{G}}\left(a,b,p,q\right)$, $N_{G}=T_{G}+2$, which determines the maximum number of iterations, $T_{G}$, for the GFS. For the $g$-th iteration, the IU is divided based on the ratio of two successive higher GFNs, i.e.,
@@
$\hat{\gamma}_{i}=\frac{{x}_{s}^{\left(g-1\right)}+{x}_{f}^{\left(g-1\right)}}{2},$ (59)
The GFS algorithm for estimating $\hat{\gamma}_{i}$ is summarized in Algorithm 2. Except for $\hat{\gamma}_{i}$, the MF-GFS scheme follows the same procedure as the MF scheme. For the MF-GFS scheme, $\hat{\gamma}_{i}$ is obtained via the Algorithm 2 rather than (50).
Note that the size of IU in the MF scheme is $\small{\frac{1}{\rho+1}}$, which is generally on the magnitude of $10^{-1}$ or $10^{-2}$. However, the size of IU in the MF-GFS scheme is several orders of magnitude smaller. Therefore, the proposed MF-GFS scheme is generally more effective than the MF scheme.
Algorithm 2 GFS Algorithm
0:  Received signal $\boldsymbol{y}_{\mathrm{T}}$ Initialization: $g=1$, $x_{s}^{\left(0\right)}=-0.5$, $x_{f}^{\left(0\right)}=0.5$, the objective function $f\left(\gamma\right)=\left|\boldsymbol{\mathcal{T}}^{\mathrm{H}}\left(\hat{\varUpsilon}_{i}+\gamma\right)\boldsymbol{y}_{\mathrm{T}}\right|^{2}$.
@@
3:  According to (55), the new IU is obtained.
@@
6:  Calculate $\hat{\gamma}_{i}$ according to (59).
6:  $\hat{\gamma}_{i}$.
Algorithm 3 The proposed MF-GFS CE scheme for IDFD channels
0:  Received signal $\boldsymbol{y}_{\mathrm{T}}$. Initialization: $\boldsymbol{y}_{\mathrm{T}}^{\left(t\right)}=\boldsymbol{y}_{\mathrm{T}}$, the maximum number of iterations $T_{\mathrm{iter}}$ and $T_{G}$, threshold $\sigma$ and $\epsilon$, $l_{\max}$, $k_{\max}$, $\hat{\boldsymbol{l}}_{\tau}=\left[\right],\hat{\boldsymbol{k}}_{\nu}=\left[\right],\hat{\boldsymbol{h}}=\left[\right]$, $t_{\mathrm{iter}}=1$.
@@
3:  Based on (61), estimate $\hat{\kappa}_{\nu}^{\left(t\right)}$ by Algorithm 2.
4:  $\hat{h}^{\left(t\right)}$ can be obtained by (46).
5:  $\hat{\boldsymbol{l}}_{\tau}=\left[\hat{\boldsymbol{l}}_{\tau},\hat{l}_{\tau}^{\left(t\right)}\right]$, $\hat{\boldsymbol{k}}_{\nu}=\left[\hat{\boldsymbol{k}}_{\nu},\hat{k}_{\nu}^{\left(t\right)}\right]$, $\hat{\boldsymbol{h}}=\left[\hat{\boldsymbol{h}},\hat{h}^{\left(t\right)}\right]$.
6:  $\boldsymbol{y}_{\mathrm{T}}^{\left(t+1\right)}$ is obtained by (54).
@@
8:  $\left(\hat{\boldsymbol{l}}_{\tau},\hat{\boldsymbol{k}}_{\nu},\hat{\boldsymbol{h}}\right)$.
IV-E MF-GFS CE scheme for IDFD channels
In typical wide-band systems such as satellite communications or millimeter-wave communications, $B$ is sufficiently large to approximate $\tau_{i}$ to the nearest sampling point [7, 38]. Therefore, an MF-GFS scheme is proposed for IDFD channels.
The observation model in (35) is still considered. The estimation of $\left(h_{i},l_{\tau_{i}}=\bar{l}_{\tau_{i}},k_{\nu_{i}}=\bar{k}_{\nu_{i}}+\kappa_{\nu_{i}}\right)$ for the $i$-th path can also be expressed as (45) and (46). Since the $\iota_{\tau_{i}}=0$, path ambiguity does not exist. Therefore, based on the pilot pattern in Fig. 3, $\left(\hat{l}_{\tau_{i}},\hat{k}_{\nu_{i}}\right)$ can be accurately estimated.
For (45), $\left(\hat{l}_{\tau_{i}},\hat{k}_{\nu_{i}}\right)$ can be obtained by searching for consecutive $\left(l_{\tau_{i}},k_{\nu_{i}}\right)$. It can be anticipated that $\left|\boldsymbol{a}_{\mathrm{T}}^{\mathrm{H}}\left(l_{\tau},k_{\nu}\right)\boldsymbol{y}_{T}\right|$ achieves its maximum value at $\left(l_{\tau},k_{\nu}\right)=\left(l_{\tau_{i}},k_{\nu_{i}}\right)$. However, the complexity of searching for $\left(\hat{l}_{\tau_{i}},\hat{k}_{\nu_{i}}\right)$ in space $\left\{\left[0,l_{\max}\right]\times\left[-k_{\max},k_{\max}\right]\right\}$ is still high. A simple and effective method is to estimate $\left(\bar{l}_{\tau_{i}},\bar{k}_{\nu_{i}}\right)$ of the path based on the maximum peaks in the received signal $\boldsymbol{y}_{\mathrm{T}}$, i.e.,
@@ IV-E MF-GFS CE scheme for IDFD channels
The estimation of $\kappa_{\nu_{i}}$ can be expressed as
@@
The MF scheme is to directly search the grid $\boldsymbol{\varGamma}$ to get $\hat{\kappa}_{\nu_{i}}$. In addition, $\left|\boldsymbol{a}_{\mathrm{T}}^{\mathrm{H}}\left(\hat{\bar{l}}_{\tau_{i}},\hat{\bar{k}}_{\nu_{i}}+\kappa_{\nu}\right)\boldsymbol{y}_{T}\right|^{2}$ is a bounded unimodal function, achieving its extreme value at $\kappa_{\nu}=\kappa_{\nu_{i}}$. This proof is similar to that of Theorem 3. Therefore, to reduce redundant computation and improve performance, Algorithm 2 is used to estimate $\hat{\kappa}_{\nu_{i}}$. Note that the same iterative method is also considered for parameter estimation of multiple paths, which estimates one path in one iteration. The proposed MF-GFS CE scheme for IDFD channels is summarized in Algorithm 3.
IV-F Complexity Analysis
In this subsection, we analyze complexity of the proposed MF and MF-GFS CE schemes. For the MF CE in the IDFD channels, The complexity of (60) is $\mathcal{O}\left(Q\right)$. In addition, based on (61), estimating $\hat{\kappa}_{\nu_{i}}$ requires $\mathcal{O}\left(Q\rho\right)$. Calculating (46) costs $\mathcal{O}\left(Q\right)$. To calculate $\boldsymbol{y}_{\mathrm{T}}^{\left(t+1\right)}$ in (54), $\mathcal{O}\left(Q\right)$ are required. Therefore, the complexity in each iteration is $\mathcal{O}\left(Q\rho\right)$. For the proposed MF-GFS CE scheme in the IDFD channel, estimating $\hat{\kappa}_{\nu_{i}}$ through Algorithm 2 costs $\mathcal{O}\left(T_{G}Q\right)$. Therefore, the complexity in each iteration is $\mathcal{O}\left(T_{G}Q\right)$. The remaining analysis is similar to the MF scheme. The complexity is summarized in Table I, where $L$ is the number of columns in the measurement matrix for the CS schemes. Fig. 8 is the complexity comparison, where $l_{\max}=8$, $k_{\max}\in\left[2,10\right]$. one can see the low-complexity nature of the proposed schemes.
@@ IV-F Complexity Analysis
The proposed MF scheme for IDFD channels $\mathcal{O}\left(T_{\mathrm{iter}}\rho Q\right)$
The proposed MF-GFS scheme for IDFD channels $\mathcal{O}\left(T_{\mathrm{iter}}T_{G}Q\right)$
The proposed MF scheme for FDFD channels $\mathcal{O}\left(T_{\mathrm{iter}}\rho Q\right)$
The proposed MF-GFS scheme for FDFD channels $\mathcal{O}\left(T_{\mathrm{iter}}T_{G}Q\right)$
@@ V Simulation Results
Data modulation $4$-QAM
$\xi$, $\xi^{\prime}$ $4$, $5$
In this section, the CE performances of the proposed MF and MF-GFS methods are evaluated. Unless otherwise stated, the simulation parameters are shown in Table II. $\tau_{i}$ and $\nu_{i}$ are randomly and uniformly generated within $\left[0,\tau_{\max}\right]$ and $\left[-\nu_{\max},\nu_{\max}\right]$, respectively. The system signal-to-noise ratio (SNR) is defined as $\frac{E\left\{\left|x\left(n\right)\right|^{2}\right\}}{\sigma^{2}}$, where $E\left\{\cdot\right\}$ is the expected operation. For the pilot $x_{p}$, $3010\log_{10}\frac{\left|x_{p}\right|^{2}}{E\left\{\left|x\left(n\right)\right|^{2}\right\}}=$ 30 dB is satisfied. For the proposed algorithms, set $T_{\mathrm{iter}}=15$, $\sigma=10^{-3}$, $\rho=20$. For comparison, we consider the traditional LS scheme [28], the MLE scheme [15], the on-grid scheme [29, 30] with OMP algorithm, and the off-grid scheme [31]. For CE schemes based on CS algorithm, the grid resolution is 0.5. We reconstruct $\hat{\boldsymbol{H}}_{\mathrm{eff}}$ using the estimated parameters, and define NMSE as $\mathrm{NMSE}\left.=\right\|\boldsymbol{H}_{\mathrm{eff}}-\hat{\boldsymbol{H}}_{\mathrm{eff}}\left\|{}^{2}\right./\left\|\boldsymbol{H}_{\mathrm{eff}}\right\|^{2}$.
Figure 9: Comparison of different AFDM frameworks.
Fig. 9 shows the BER performance comparison for different AFDM frameworks, whereby the BER is obtained by LMMSE algorithm. The performances of the two AFDM frameworks are consistent. In addition, if the coupling phase $e^{\small{\frac{-j2\pi}{N}k_{\nu_{i}}l_{\tau_{i}}}}$ is not considered, the BER will be severely deteriorated.
@@
The approximate orthogonality of $\boldsymbol{A}_{\mathrm{T}}\left(\boldsymbol{l}_{\tau},\boldsymbol{k}_{\nu}\right)$ for different values of $\xi$ and $N$ is shown in Fig. (10). For “$\text{NMSE}_{\text{A}}$ vs. $\xi$”, we consider $N=256$, $\xi\in\left[1,8\right]$. For “$\text{NMSE}_{\text{A}}$ vs. $N$”, we have $\xi=4$, $N\in 200\times\left[1,8\right]$. To make this orthogonality more obvious, the integer parts of the delays for different paths are set to differ. Note that $\text{NMSE}_{\text{A}}\left.=\right\|\boldsymbol{I}-\boldsymbol{A}_{\mathrm{T}}^{\mathrm{H}}\left(\boldsymbol{l}_{\tau},\boldsymbol{k}_{\nu}\right)\boldsymbol{A}_{\mathrm{T}}\left(\boldsymbol{l}_{\tau},\boldsymbol{k}_{\nu}\right)\left\|{}^{2}\right./\left\|\boldsymbol{I}\right\|^{2}$. It can be observed that this approximate orthogonality is still good enough even in fractional channels. This orthogonality improves with the increase of $\xi$ or $N$ in both IDFD and FDFD channels, which is also confirmed in Appendix B.
Figure 11: NMSE comparison of the proposed MF scheme versus $\rho$ for different SNR values.
@@
Fig. 11 and 12 show the NMSE and BER performances of the proposed MF scheme for IDFD channel with different SNR and $\rho$, respectively. As $\rho$ increases, the size of IU correspondingly decreases, enabling more accurate estimation of the fractional Doppler, which leads to improvements in both NMSE and BER.
@@
The impact of different $\xi$ on the BER performance of the MF scheme is shown in Fig. 13. It can be observed that the BER performance improves as the $\xi$ increases. In this case, the interference between different paths will decrease, making the estimated parameters more accurate, and hence improved BER performance. In addition, according to Appendix B and Fig. 10, as $\xi$ increases, the orthogonality performance of $\boldsymbol{A}_{\mathrm{T}}\left(\boldsymbol{l}_{\tau},\boldsymbol{k}_{\nu}\right)$ is improved, thus leading to further improved CE and BER performances.
@@
The NMSE and BER performances of different schemes for IDFD channels are compared in Fig. 14 and Fig. 15, respectively. We set $\rho=10$ and $N_{G}=15$ to allow a comparison of the MF scheme and the MF-GFS scheme at almost the same complexity levels. In addition, the performance of the MF-GFS scheme with different GFN, distinguished by $\left(a,b,p,q\right)$, is also compared. It can be observed that the performance of the proposed MF scheme and MF-GFS scheme are significantly improved compared to that of other schemes. Since the size of the IU estimated by the MF-GFS scheme is significantly smaller than that of MF, it can be seen that the performance of the proposed MF-GFS scheme outperforms that of the MF scheme at high SNR. In addition, since the ratio of two consecutive GFNs is generally close to the golden ratio, different GFNs yield similar IU divisions, thereby resulting in a negligible impact on the performance of the MF-GFS scheme.
Fig. 16 is a comparison of BER performance for FDFD channels. For simple verification, assume that the relative positions of the peaks of the pilot responses in $\boldsymbol{y}_{\mathrm{T}_{1}}$ and $\boldsymbol{y}_{\mathrm{T}_{2}}$ are known. Since the LS, MLE, on-grid and off-grid schemes fail to address the path ambiguity caused by fractional delay, they suffer from poor BER performance. In contrast, the proposed MF and MF-GFS schemes outperform significantly as they can achieve accurate estimation in FDFD channels.

### v3

@@ IV-B MF CE for FDFD channels
2: According to (50), $\left(\bar{l}_{\tau}^{\left(t\right)},\bar{k}_{\nu}^{\left(t\right)}\right)$ is obtained.
[…]
6: The $\boldsymbol{y}_{\mathrm{T}}^{\left(t+1\right)}$ is obtained by (55).
7:$t=t+1$.
[…]
IV-C MF-GFS-DE CE for FDFD channels
In Subsection IV-B, an MF CE scheme is proposed. However, the trade-off between the complexity and estimation performance of the MF CE scheme yet depends on the $\rho$. Performing a global search over an interval is inefficient, and the estimation accuracy is limited by the search step size $\rho$. The larger the $\rho$, the finer the grid $\boldsymbol{\varGamma}$, enabling more accurate estimation of fractional part at the cost of higher complexity. To address this issue, we introduce a MF-GFS CE scheme based on the GFS algorithm in this subsection. The GFS algorithm unevenly divides the search interval into two parts through the ratio of two consecutive generalized Fibonacci numbers (GFNs) to achieve an efficient search. The $i$-th GFN is denoted by $S_{i}\left(a,b,p,q\right)$, where $S_{0}=a$, $S_{1}=b$, $S_{i+2}=pS_{i+1}+qS_{i}$, $i$ is a non-negative integer, $a$ and $b$ be two non-negative integers such that $a+b>0$, $p$ and $q$ be two positive integers.
Note that an important pre-condition for using the GFS algorithm is that the objective function should be a bounded unimodal function. Since fractional delays may cause local peaks, the objective functions in both (51) and (52) are generally not unimodal.
@@ IV-C MF-GFS-DE CE for FDFD channels
In (53), $\left|\boldsymbol{a}_{\mathrm{T}}^{\mathrm{H}}\left(\hat{\bar{l}}_{\tau_{i}}+\hat{\iota}_{\tau_{i}},\hat{\bar{k}}_{\nu_{i}}+\kappa_{\nu}\right)\boldsymbol{y}_{T}\right|^{2}$ is a bounded unimodal function.
@@
After dividing the search interval by two consecutive GFNs, the interval of uncertainty (IU) containing the optimal value $\kappa_{\nu_{i}}$ needs to be determined. For convenience, assume that the search interval is $\left[x_{s},x_{f}\right]$, and $x_{s}<x_{1}<x_{2}<x_{f}$. The new IU $\left[x_{s}^{\prime},x_{f}^{\prime}\right]$ can be expressed as
@@
The process of the GFS for estimating $\hat{\kappa}_{\nu_{i}}$ is shown in Fig. 4. Note that it is necessary to set a maximum GFN $S_{N_{G}}$, $N_{G}=T_{G}+2$, which determines the maximum number of iterations, $T_{G}$, for the GFS. For the $g$-th iteration, the IU is divided based on the ratio of two successive higher GFNs, i.e.,
@@
$\hat{\kappa}_{\nu_{i}}=\frac{{x}_{s}^{\left(g-1\right)}+{x}_{f}^{\left(g-1\right)}}{2}.$ (60)
The proposed GFS algorithm for estimating $\hat{\kappa}_{\nu_{i}}$ is summarized in Algorithm 2. Except for $\hat{\kappa}_{\nu_{i}}$, the MF-GFS scheme follows the same procedure as the MF scheme. For the MF-GFS scheme, $\hat{\kappa}_{\nu_{i}}$ is obtained via the Algorithm 2 rather than (53).
Note that the size of IU in the global search method is $\small{\frac{1}{\rho+1}}$, which is generally on the magnitude of $10^{-1}$ or $10^{-2}$. However, the size of IU in the GFS is several orders of magnitude smaller. Therefore, the proposed MF-GFS scheme is generally more effective than the MF scheme.
Algorithm 2 The proposed GFS Algorithm
0: Received signal $\boldsymbol{y}_{\mathrm{T}}$, the objective function $f\left(\kappa_{\nu}\right)=\left|\boldsymbol{a}_{\mathrm{T}}^{\mathrm{H}}\left(\hat{\bar{l}}_{\tau_{i}}+\hat{\iota}_{\tau_{i}},\hat{\bar{k}}_{\nu_{i}}+\kappa_{\nu}\right)\boldsymbol{y}_{T}\right|^{2}$. Initialization: $g=1$, $T_{G}$, IU $\left[x_{s}^{\left(0\right)},x_{f}^{\left(0\right)}\right]=\left[-0.5,0.5\right]$.
@@
3: According to (56), the new IU $\left[x_{s}^{\left(g\right)},x_{f}^{\left(g\right)}\right]$ is obtained.
@@
6: Calculate $\hat{\kappa}_{\nu_{i}}$ according to (60).
[…]
Algorithm 3 The proposed MF and MF-GFS CE scheme for IDFD channels
[…]
@@
[…]
@@
IV-D MF and MF-GFS CE schemes for IDFD channels
In typical wide-band systems such as satellite or millimeter-wave communications, $B$ is sufficiently large to approximate $\tau_{i}$ to the nearest sampling point [6]. Therefore, MF and MF-GFS schemes are proposed for IDFD channels.
The observation model in (40) is still considered. The estimation of $\left(h_{i},l_{\tau_{i}}=\bar{l}_{\tau_{i}},k_{\nu_{i}}=\bar{k}_{\nu_{i}}+\kappa_{\nu_{i}}\right)$ for the $i$-th path can also be expressed as (48) and (49).
To reduce complexity, the integer $\left(\bar{l}_{\tau_{i}},\bar{k}_{\nu_{i}}\right)$ is initially also estimated using (50). $\hat{\kappa}_{\nu_{i}}$ can be expressed as
@@ IV-D MF and MF-GFS CE schemes for IDFD channels
The MF scheme is to directly search the grid $\boldsymbol{\varGamma}$ to get $\hat{\kappa}_{\nu_{i}}$. In addition, $\left|\boldsymbol{a}_{\mathrm{T}}^{\mathrm{H}}\left(\hat{\bar{l}}_{\tau_{i}},\hat{\bar{k}}_{\nu_{i}}+\kappa_{\nu}\right)\boldsymbol{y}_{T}\right|^{2}$ is a bounded unimodal function, achieving its extreme value at $\kappa_{\nu}=\kappa_{\nu_{i}}$. This proof is similar to Appendix C. Therefore, to reduce complexity and improve performance, Algorithm 2 is used to estimate $\hat{\kappa}_{\nu_{i}}$. Note that the same iterative method is also considered for parameter estimation of multiple paths, which estimates one path in one iteration. The proposed MF and MF-GFS CE schemes for IDFD channels are summarized in Algorithm 3.
IV-E Complexity Analysis
In this subsection, we analyze complexity of the proposed MF and MF-GFS CE schemes. For the MF CE in the IDFD channels, The complexity of (50) is $\mathcal{O}\left(Q\right)$. In addition, based on (61), estimating $\hat{\kappa}_{\nu_{i}}$ requires $\mathcal{O}\left(Q\rho\right)$. Calculating (49) costs $\mathcal{O}\left(Q\right)$. To calculate $\boldsymbol{y}_{\mathrm{T}}^{\left(t+1\right)}$ in (55), $\mathcal{O}\left(Q\right)$ are required. Therefore, the complexity in each iteration is $\mathcal{O}\left(Q\rho+3Q\right)$. For the proposed MF-GFS CE scheme in the IDFD channel, estimating $\hat{\kappa}_{\nu_{i}}$ through Algorithm 2 costs $\mathcal{O}\left(T_{G}Q\right)$. Therefore, the complexity in each iteration is $\mathcal{O}\left(T_{G}Q+3Q\right)$. The remaining analysis is similar to the MF scheme. The complexity is summarized in Table I, where $L$ is the number of columns in the measurement matrix for the CS schemes. Fig. 5 is the complexity comparison, where $l_{\max}=8$, $k_{\max}\in\left[2,10\right]$. one can see the low-complexity nature of the proposed schemes.
@@ IV-E Complexity Analysis
The proposed MF scheme for IDFD channels $\mathcal{O}\left(T_{\mathrm{iter}}\left(\rho Q+3Q\right)\right)$
The proposed MF-GFS scheme for IDFD channels $\mathcal{O}\left(T_{\mathrm{iter}}\left(T_{G}Q+3Q\right)\right)$
The proposed MF-JE and MF-DE scheme for FDFD channels $\mathcal{O}\left(T_{\mathrm{iter}}\left(\rho^{2}Q+3Q\right)\right)$, $\mathcal{O}\left(T_{\mathrm{iter}}\left(2\rho Q+3Q\right)\right)$
The proposed MF-GFS-DE scheme for FDFD channels $\mathcal{O}\left(T_{\mathrm{iter}}\left(\rho+T_{G}\right)Q+3Q\right)$
@@ V Simulation Results
Data modulation 44-QAM
$\xi$, $\xi^{\prime}$ 44, 55
In this section, the CE performances of the proposed MF and MF-GFS methods are evaluated. Unless otherwise stated, the simulation parameters are shown in Table II. $\tau_{i}$ and $\nu_{i}$ are randomly and uniformly generated within $\left[0,\tau_{\max}\right]$ and $\left[-\nu_{\max},\nu_{\max}\right]$, respectively. The system signal-to-noise ratio (SNR) is defined as $\frac{E\left\{\left|x\left[n\right]\right|^{2}\right\}}{\sigma^{2}}$. For the pilot $x_{p}$, $10\log_{10}\frac{\left|x_{p}\right|^{2}}{E\left\{\left|x\left[n\right]\right|^{2}\right\}}=30$ dB is satisfied. For the proposed algorithms, set $T_{\mathrm{iter}}=15$, $\sigma=10^{-3}$, $\rho=20$. For comparison, we consider the traditional LS scheme [37], the MLE scheme [16], the on-grid scheme [38, 39] with OMP algorithm, and the off-grid scheme [40]. For CE schemes based on CS algorithm, the grid resolution is 0.5. We reconstruct $\hat{\boldsymbol{H}}_{\mathrm{eff}}$ using the estimated parameters, and define NMSE as $\mathrm{NMSE}\left.=\right\|\boldsymbol{H}_{\mathrm{eff}}-\hat{\boldsymbol{H}}_{\mathrm{eff}}\left\|{}^{2}\right./\left\|\boldsymbol{H}_{\mathrm{eff}}\right\|^{2}$.
Figure 6: Complexity of constructing $\boldsymbol{H}_{\mathrm{eff}}$ via different methods.
Fig. 6 shows a comparison of the complexity of constructing $\boldsymbol{H}_{\mathrm{eff}}$ using different methods. For large $N$, the proposed FFT-based and SS methods reduce complexity by at least two orders of magnitude. The SS method is preferable for small $R$, while the FFT-based method is better for large $R$.
@@
The approximate orthogonality of $\boldsymbol{A}_{\mathrm{T}}\left(\boldsymbol{l}_{\tau},\boldsymbol{k}_{\nu}\right)$ for different values of $\xi$ and $N$ is shown in Fig. (7). For $\text{NMSE}_{\text{A}}$ vs. $\xi$”, we consider $N=256$, $\xi\in\left[1,8\right]$. For $\text{NMSE}_{\text{A}}$ vs. NN”, we have $\xi=4$, $N\in 200\times\left[1,8\right]$. To make this orthogonality more obvious, the integer parts of the delays for different paths are set to differ. Note that $\text{NMSE}_{\text{A}}\left.=\right\|\boldsymbol{I}-\boldsymbol{A}_{\mathrm{T}}^{\mathrm{H}}\left(\boldsymbol{l}_{\tau},\boldsymbol{k}_{\nu}\right)\boldsymbol{A}_{\mathrm{T}}\left(\boldsymbol{l}_{\tau},\boldsymbol{k}_{\nu}\right)\left\|{}^{2}\right./\left\|\boldsymbol{I}\right\|^{2}$. It can be observed that this approximate orthogonality is still good enough even in fractional channels. This orthogonality improves with the increase of $\xi$ or $N$. However, for FDFD channels, increasing $\xi$ does not improve orthogonality.
Figure 8: Effects of $\rho$ on MF CE and $T_{G}$ on MF-GFS CE.
@@
Fig. 8 shows the effects of $\rho$ on MF and $T_{G}$ on MF-GFS CE for IDFD channel. Compared to the MF CE scheme, the MF-GFS CE scheme can achieve better performance with lower complexity. Fig. 9 show the BER performance of the proposed MF scheme for IDFD channel with different SNR and $\rho$, respectively. The BER is obtained by LMMSE algorithm. As $\rho$ increases, the size of IU correspondingly decreases, enabling more accurate estimation of the fractional Doppler, which leads to improvements in both NMSE and BER.
@@
The impact of different $\xi$ on the BER of the MF scheme is shown in Fig. 10. It can be observed that the BER improves as the $\xi$ increases. In this case, the interference between different paths will decrease, making the estimated parameters more accurate, and hence improved BER performance. In addition, according to Appendix B and Fig. 7, as $\xi$ increases, the orthogonality performance of $\boldsymbol{A}_{\mathrm{T}}\left(\boldsymbol{l}_{\tau},\boldsymbol{k}_{\nu}\right)$ is improved, thus leading to further improved CE and BER performances.
@@
The NMSE and BER performances of different schemes for IDFD channels are compared in Fig. 11 and Fig. 12, respectively. We set $\rho=15$ and $T_{G}=8$. In addition, the performance of the MF-GFS scheme with different GFN, distinguished by $\left(a,b,p,q\right)$, is also compared. The performance of the proposed MF scheme and MF-GFS scheme are significantly improved compared to that of other schemes. Since the size of the IU estimated by the MF-GFS scheme is significantly smaller than that of MF, it can be seen that the performance of the proposed MF-GFS scheme outperforms that of the MF scheme at high SNR. In addition, since the ratio of two consecutive GFNs is generally close to the golden ratio, different GFNs yield similar IU divisions, thereby resulting in a negligible impact on the performance of the MF-GFS scheme.
Fig. 13 is a comparison of BER performance for FDFD channels. As analyzed in subsection III-B-3), local peaks induced by $\iota_{\tau_{i}}$ exacerbate ICCI and inter-path interference, leading to deteriorated CE and BER performances. However, the proposed MF CE scheme outperforms existing schemes.